OpenStudy (anonymous):

The functions f(x) and g(x) are described using the following equation and table: f(x) = 3(1.02)^x x g(x) -1 -4 0 6 1 8 2 10 Which statement best compares the y-intercepts of f(x) and g(x)? The y-intercept of f(x) is equal to the y-intercept of g(x). The y-intercept of f(x) is equal to 2 times the y-intercept of g(x). The y-intercept of g(x) is equal to 2 times the y-intercept of f(x). The y-intercept of g(x) is equal to 2 plus the y-intercept of f(x).

OpenStudy (kirbykirby):

The y-intercept is where x=0, so solve for x=0 in f(x) \(f(x) = 3(1.02)^0=3(1)=3\) for g(x), they give you a table and directly tell you that for x=0, g(x) 6 so the intercept for g(x) is 6

OpenStudy (anonymous):

@kirbykirby : so which awnser is it? The y-intercept of f(x) is equal to the y-intercept of g(x). The y-intercept of f(x) is equal to 2 times the y-intercept of g(x). The y-intercept of g(x) is equal to 2 times the y-intercept of f(x). The y-intercept of g(x) is equal to 2 plus the y-intercept of f(x).

OpenStudy (kirbykirby):

for f(x) it's 3 for g(x) it's 6 how to you get from 3 to 6 according to the choices given?

OpenStudy (anonymous):

c? @kirbykirby

OpenStudy (kirbykirby):

yes indeed :)!

OpenStudy (anonymous):

Yay! do you mind helping me some other questions?

OpenStudy (kirbykirby):

ok sure

OpenStudy (anonymous):

@kirbykirby

OpenStudy (anonymous):

The graph of f(x) =(0.5)^x is replaced by the graph of g(x) =(0.5)^x-k. If g(x) is obtained by shifting f(x) down by 5 units, the value of k is _________.

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

5, since there is already a negative in front of k

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

Which function do you think could represent f(x), and which one do you think represents g(x) ?

OpenStudy (anonymous):

B?

OpenStudy (kirbykirby):

not quite

OpenStudy (kirbykirby):

if \(f(x)=4^x\) and \(g(x)=-1\) then \(f(x) + g(x) = 4^x + (-1)\)

OpenStudy (anonymous):

oh so D

OpenStudy (kirbykirby):

yup

OpenStudy (anonymous):

Sorry this is part of my finals @kirbykirby

OpenStudy (anonymous):

would you mind helping me with the questions?

OpenStudy (kirbykirby):

are there a lot more

OpenStudy (anonymous):

@kirbykirby

OpenStudy (anonymous):

5 more

OpenStudy (anonymous):

actually help me with these two and ill be gone lol @kirbykirby :)

OpenStudy (kirbykirby):

11, does it look exponential or linear?

OpenStudy (anonymous):

linear

OpenStudy (kirbykirby):

if it's linear, then the amount of increase will be the same for x=1, x=2, x=3... So like if x=1 with value 50 x=2 with value 60 x=3 with value 70 this is a linear increase since it's always increasing by the same amount 10.

OpenStudy (kirbykirby):

if you subtract the values from x=3 and x=2, and compare it to the subtraction of x=2 and x=1 and see that its not the same, it will be exponential, if it's the same then it's linear

OpenStudy (anonymous):

C!

OpenStudy (kirbykirby):

it would actually be D. The equal increase by a factor would imply a multiplication, whereas an increase in number would mean a simple increase as I mentioned above (+10, +10...) The function represented would actually be \(\large 50(5)^{x-1}\), so it's constantly increasing by a factor of 5 (meaning you constantly multiply that number by 5 50 -> 250, multiply by 5 250 -> 1250 , multiply by 5

OpenStudy (anonymous):

Ohhhhhhhhh ok i got it :) thanks and what about 12?

OpenStudy (kirbykirby):

Try to determine which options are increasing by adding a number every time, or by multiplying a number every time If you are adding : it's linear If you are multiplying: it's exponential

OpenStudy (anonymous):

so C?

OpenStudy (kirbykirby):

in option one, you keep doing +0.3, +0.3.. so it can't be exponential

OpenStudy (anonymous):

Ohhh okay :) thank you very much!

OpenStudy (kirbykirby):

=]